INFRARED ABSORPTION OF GASEOUS c-CLCOOH AND t-CLCOOH RECORDED WITH A STEP-SCAN FOURIER-TRANSFORM SPECTROMETER

\underline{YUAN-PERN LEE} AND LI-KANG CHU, Department of Applied Chemistry and Institute of Molecular Science, National Chiao Tung University, Hsinchu 30010, Taiwan and Institute of Atomic and Molecular Sciences, Academia Sinica, Taipei 10617, Taiwan.

Two conformers of ClCOOH were produced upon irradiation at 355 nm of a gaseous flowing mixture of Cl_{2}, HCOOH, and N_{2}. A step-scan Fourier-transform infrared spectrometer coupled with a multipass absorption cell was utilized to monitor the transient spectra of ClCOOH. Absorption bands with origins at 1808.0 and 1328.5 cm^{−1} are attributed to the C=O stretching and COH bending modes of t-ClCOOH, respectively; those at 1883.0 and 1284.9 cm^{−1} are assigned as the C=O stretching and COH bending modes of c-ClCOOH, respectively. These observed vibrational wavenumbers agree with corresponding values for t-ClCOOH and c-ClCOOH predicted with B3LYP/aug-cc-pVTZ density-functional theory and the observed rotational contours agree satisfactorily with simulated bands based on predicted rotational parameters. The observed relative intensities indicate that t-ClCOOH is more stable than c-ClCOOH by 3 kJ mol^{−1}. A simple kinetic model is employed to account for the production and decay of ClCOOH.